OpenStudy (anonymous):

Someone help please i dont know how to work this problem.. Solve. Round your answer to three decimal places. 5^x+1=4^x-1

OpenStudy (anonymous):

i am going to assume that this is \[5^{x+1}=4^{x-1}\]

OpenStudy (anonymous):

satellite always gives the best answers

OpenStudy (anonymous):

@satellite73 yes

OpenStudy (anonymous):

start with \[(x+1)\ln(5)=(x-1)\ln(4)\] and then do a raft of algebra is the first step clear?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then it is algebra from here on in don't forget that while \(\ln(x)\) is a function, \(\ln(5)\) is a number, some constant

OpenStudy (anonymous):

-13.425 :)

OpenStudy (anonymous):

so we start with \[(x+1)\ln(5)=(x-1)\ln(4)\] multiply out to get \[\ln(5)x+\ln(5)=\ln(4)x-\ln(4)\] then put the x's on one side via \[\ln(5)x-\ln(4)x=-\ln(4)-\ln(5)\] factor out the x to get \[(\ln(5)-\ln(4))x=-\ln(4)-\ln(5)\] then divide to get \[x=\frac{-\ln(4)-\ln(5)}{\ln(5)+\ln(4)}\]